OpenStudy (anonymous):

Which number sentence shows the associative property of addition? Choose all that apply. 3z + 9p = 9p + 3z 3x (5 + 2) = 3x (5) + 3x (2) 3 + (2 + 8) = (3 + 2) + 8 (6 + 1) + 9 = 6 + (1 + 9)

OpenStudy (shamil98):

third one..

OpenStudy (anonymous):

3 + (2 + 8) = (3 + 2) + 8 (6 + 1) + 9 = 6 + (1 + 9)

OpenStudy (shamil98):

Read this: http://www.mathsisfun.com/associative-commutative-distributive.html

OpenStudy (anonymous):

Thanks @shamil98 and @wangkan07

OpenStudy (anonymous):

medal?